Problem 6.
Complete a common-size vertical analysis in the condensed statement of comprehensive income presented below:

Condensed Statement of Comprehensive Income
Sales revenue
P 480,000

Cost of sales
201,600

Gross margin
278,400

Operating expenses
206,400

Operating income
P 72,000
1
Expert's answer
2021-09-28T17:42:27-0400

Common-size Vertical analysis = statement line item/Total base figure × 100

Total base figure in income statement is usually gross sales.


Sales revenue

P 480,000 =480,000/480,000 ×100 =100%

Cost of sales

201,600 = 201,600/480,000 × 100 = 42%

Gross margin

278,400 = 278,400/480,000 × 100 = 58%

Operating expenses

206,400 = 206,400/480,000 × 100 = 43%

Operating income

P 72,000 = 72,000/480,000 × 100 = 15%